OpenStudy (vijeya3):

What is the remainder when 2^33 is divided by 17?

OpenStudy (nubeer):

maybe first find 2^33

OpenStudy (vijeya3):

really? you think i should do that tedious work? it can't be that straight..

ganeshie8 (ganeshie8):

what else have you tried ?

OpenStudy (nubeer):

i know .. its more like some theorem

Parth (parthkohli):

aren't you too young to be familiar with modular arithmetic and/or binomial theorem?

OpenStudy (vijeya3):

@parth I dunno it is from my textbook:(

Parth (parthkohli):

so then you're not familiar with any of those?

OpenStudy (vijeya3):

@ganeshie8 I tried the cyclicity but that doesn't give any clue at all..... @ParthKohli What ya mean?

Parth (parthkohli):

binomial theorem modular arithmetic do you know any of these? and yes, the cyclic nature of remainders should give a nice hint

ganeshie8 (ganeshie8):

thats not a bad attempt! familiar with below property ? If \(a\) leaves a remainder \(r\) when divided by \(b\), then \(a^k\) leaves a remainder \(r^k\) when divided by \(b\)

ganeshie8 (ganeshie8):

for example : \(5\) leaves a remainder \(1\) when divided by \(4\), therefore \(5^3\) leaves a remainder \(1^3\) when divided by \(4\)

ganeshie8 (ganeshie8):

you might need something like binomial theorem to prove above property

OpenStudy (vijeya3):

@ParthKohli I am not familiar with any of those:( and @ganeshie8 I dunno that property but yeah..that gives some insight...

OpenStudy (mathmath333):

u shuold study negative remainders then this questions will be easy

OpenStudy (vijeya3):

negative remainders? okay @ganeshie8 so if 2 leaves a remainder 15 when divided by 17,therefore 2^33 leaves a remainder 15^33 when divided by 17???:/

OpenStudy (vijeya3):

@ParthKohli what hint from cyclicity please?

Parth (parthkohli):

observe that \(2^4\) leaves a remainder \(-1\) when divided by 17. or better, \(2^8\) leaves a remainder \(1\) when divided by 17. whenever you get a remainder of 1, the remainders begin to repeat themselves.

OpenStudy (vijeya3):

@ParthKohli 'whenever you get a remainder of 1, the remainders begin to repeat themselves.' What does this mean?I dun understand:(

Parth (parthkohli):

a simpler example would be this: remainders of powers of 2 when divided by 5. 2^1: 2 2^2: 4 2^3: 3 2^4: 1 2^5: 2 2^6: 4 2^7: 3 2^8: 1 do you see how it's going like 2, 4, 3, 1, 2, 4, 3, 1...? and the last number of every cycle is always 1.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &\dfrac{2^{33}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2^{33}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2\cdot 2^{32}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2\cdot 4^{16}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2\cdot 16^{8}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2\cdot (-1)^{8}}{17}\hspace{.33em}\\~\\ &\implies \dfrac{2\cdot 1}{17}\hspace{.33em}\\~\\ &\implies 2\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (samigupta8):

2^33 can be written as 2(16)^8 and that can be simplified as 2(17-1)^8 now apply binomial u wil get it..

OpenStudy (vijeya3):

Well @mathmath I dun understand the last step..the one where you replaces 16^8 with (-1)^8 since I dun know what ever method you have applied there.I used the clue given by @ParthKohli in order to procure the answer and got 2 that is the same as yours..but my textbook says 15 which is really weird since 2 seems correct....

OpenStudy (vijeya3):

@samigupta8 I am not familiar with the binomial theorem...but thanks for lending a hand here:)

OpenStudy (mathmath333):

if (a-1) is divided by a where a>1 and is a natural numberthen remainder can be written as -1

OpenStudy (samigupta8):

Ok..no problem..well the most apt way is then as given by parth so u should follow the multiplicity part only..and 2 is right ans ..

OpenStudy (vijeya3):

Okay...so I got it!Maybe it is a misprint in my textbook Thanks everyone:) and @ParthKohli thanks for the simple example..made it much easier to understand:)

OpenStudy (mayankdevnani):

I think Fermat's little theorem is good for this problem

OpenStudy (mayankdevnani):

@Vijeya3 https://en.wikipedia.org/wiki/Fermat's_little_theorem

OpenStudy (vijeya3):

@mayankdevnani honestly?I couldn't make head or tail of that theorem.. Anyways,thanks for lending a hand here

OpenStudy (mayankdevnani):

abe dafod ! main kisliye hoon pgir

OpenStudy (mayankdevnani):

*phir

OpenStudy (mayankdevnani):

chambal main kudd gayi kya ?

OpenStudy (vijeya3):

:( main duffer kyu? wese pata tha...phele se hi

OpenStudy (mayankdevnani):

maine kab kaha tum daffar hoon ? arrey yeh toh mazak tha bulane ke liye sorry agar tumhe bura laga ho toh

OpenStudy (vijeya3):

LOL wow!ese hi bulaya karo..kam se kam sun toh leti hu:P and yeah..i didn't get offended..sunne ki aadat h:P

OpenStudy (mayankdevnani):

let's start

OpenStudy (vijeya3):

hmm..okay warning-and if i don't understand ,dun get annoyed..m hard at grasping

OpenStudy (mayankdevnani):

pehle wikipedia pr doosra formula pado

OpenStudy (vijeya3):

dekho...wo statement samjh me aa gaya..but i dun understand that mod and all

OpenStudy (mayankdevnani):

mod aise hi cheez hain...faltu ki

OpenStudy (vijeya3):

LOL then wut is the significance of that?

OpenStudy (mayankdevnani):

baad main pata chal jayega

OpenStudy (mayankdevnani):

ab dekho, p=17 lo kyunki mod ke peeche jo bhi cheez aayegi woh divisior hogs

OpenStudy (mayankdevnani):

*hoga

Parth (parthkohli):

if she knew modular arithmetic, then she wouldn't need to know fermat's little theorem...

OpenStudy (mayankdevnani):

there are 3 methods :- 1)Binomial Theorem 2)Cyclicity 3) Fermat's little theorem

OpenStudy (mayankdevnani):

she doesn't know 1st. She knows 2nd and i teach her 3rd

Parth (parthkohli):

ek baar solve kariyo using FLT?

OpenStudy (mayankdevnani):

haan kyun nahi

OpenStudy (mayankdevnani):

let p=17 apply prime no divisibility method, \[\large \bf 2^{17-1}=1 \mod 17\] \[\large \bf 2^{16}=1 \mod 17\] since 32 is perfectly divisible by 32, so \[\large \bf 2^{32}=1 \mod 17\] multiply both sides by 2, \[\large \bf 2^{33}=2 \mod 17\] Hence, 2 is remainder

OpenStudy (mayankdevnani):

typo correction:- since 32 is perfectly divisible by 16

Parth (parthkohli):

you applied many more properties of modular arithmetic. i think if we just teach her those properties of modular arithmetic, then it'd be easier to solve. FLT is just an added burden.\[2^{33} \equiv 2\cdot(2^4)^8 \equiv 2\cdot(-1)^8 \equiv 2 \pmod{17}\]

OpenStudy (mayankdevnani):

cyclicity is best for her level now

Parth (parthkohli):

aur kya -_-

OpenStudy (vijeya3):

-_- wow

OpenStudy (mathmath333):

vijeya try these 2^31 is divided by 17

OpenStudy (vijeya3):

@mathmath333 9?

OpenStudy (mathmath333):

how u get there